QUESTION 6A business is evaluating a project for which the following information is relevant:Sales will be K100,000 in the first year and are expected to increase by 5% per year.Costs will be K50,000 and are expected to increase by 7% per year.Capital investment will be K200,000 and attracts tax allowable depreciation of the full value of the investment over the 5 year length of the project.The tax rate is 30%.The business uses a real discount rate of 9%.Calculate the NPV for the project. (15 marks)It is generally argued that the corporate objective of shareholder wealth maximization should override that of profit maximization. Distinguish these objectives and explain why wealth maximization should take precedence over profit maximization. (10 marks)(Total: 25 marks)
1
Expert's answer
2021-08-11T09:17:58-0400

"NPV=-200000+\frac{100000-50000-(100000-50000-40000)*0.3}{1.09}+\frac{100000*1.05-50000*1.07-(100000*1.05-50000*1.07-40000)*0.3}{1.09^2}+\frac{100000*1.05^2-50000*1.07^2-(100000*1.05^2-50000*1.07^2-40000)*0.3}{1.09^3}+\frac{100000*1.05^3-50000*1.07^3-(100000*1.05^3-50000*1.07^3-40000)*0.3}{1.09^4}++\frac{100000*1.05^4-50000*1.07^4-(100000*1.05^4-50000*1.07^4-40000)*0.3}{1.09^5}=-27474.33"

Shareholders are the owners of the business. In addition to dividends, they earn on the growth in the value of their asset, that is, the company in the form of shares. Therefore, profit is not the main goal, as the market capitalization of the company must also grow. Well, the value of the company depends on many factors other than profit, which ultimately affect the value of the company's shares.